OpenStudy (anonymous):

Evaluate 23 - 7 ÷ (1/3 + 1/4) show work.

OpenStudy (mattfeury):

Order of operations!! Start with the parentheses, then do the the division, then you can subtract.

OpenStudy (anonymous):

23-7 / 4+3/12 16 *12/7 192/7

OpenStudy (mattfeury):

\[\frac{1}{3} + \frac{1}{4} = \frac{4}{12} + \frac{3}{12} = \frac{7}{12}\]

OpenStudy (mattfeury):

\[23 - 7 \div \frac{7}{12} = 23 - 7 * \frac{12}{7} = 11\]